# Vacuum Polarization, Geodesic Equation and Sachs-Wolfe Effect

### v1

@@ Abstract
A key issue is now to determine $\phi^{*}\phi\right>_{L}$ and $\phi^{*}\phi\right>_{0}$. The standard dimensional regularization implies $\phi^{*}\phi\right>\sim m^{2}$, however for $m\sim$ TeV, which is a reasonably small value, $\mu\gg 1$ unless $e$ is extremely small (the current bound on $\mu$ is $|\mu|<9\times 10^{-5}$ from COBE/FIRAS and proposed experiments like PIXIE aim to to push this bound to $|\mu|\sim 10^{-8}$). A large $\mu$ appears from the same root cause as the cosmological constant problem, i.e. there is a fundamental lack of our understanding about the cosmological impact of coincident vacuum expectation values.
One can try to estimate $\phi^{*}\phi\right>$ by applying the same amount of mismatch encountered in the cosmological constant problem. The naive dimensional regularization gives $\rho\right>\sim m^{4}$, where $\rho$ is the energy density, and assuming this must have the same order of magnitude as the dark energy today gives the rough mismatch factor $H_{0}^{2}M_{p}^{2}/m^{4}$, where $H_{0}$ is the current Hubble value. Because $\rho$ has mass dimension 4 and $\phi^{*}\phi$ has mass dimension 2, the mismatch for $\phi^{*}\phi\right>_{0}$ can be $H_{0}M_{p}/m^{2}$, which implies $\phi^{*}\phi\right>_{0}\sim H_{0}M_{p}$. Taking $\omega_{0}\sim T_{0}\sim 10^{-4}eV$, the second term in (26) gives $\mu\sim 10^{-5}e^{2}$, which can be within the reach of the upcoming measurements if $e\sim 1/10$.

### v2

@@ Abstract
It is crucial to emphasize that the above analysis does not imply photons acquire mass at the fundamental level. Eq. (10) is an effective mean description of photon trajectories. This is similar to light propagation in a medium where the photons does not acquire a mass but still their speed in the medium is less than the speed of light in vacuum. In such cases, photons scatter off charged particles making their average path timelike, see FIG 1.
Figure 1: An illustration of photon propagation from point A to point B in two-dimensional spacetime. The actual path is (piece-wise) null (between collisions) while the mean path becomes timelike.
@@
A key issue is now to determine $\phi^{*}\phi\right>_{L}$ and $\phi^{*}\phi\right>_{0}$. We estimate these quantities within the free theory—neglecting any loop contributions from interactions—by assuming that the scalar field was initially in its Bunch-Davies vacuum during inflation in the early universe. Naturally, these expectation values exhibit UV divergences since the field operators have coincident spacetime coordinates and thus they must be regularized.
A standard method is to use dimensional regularization; however, it is difficult to apply in this context, as the analytical form of the Bunch-Davies mode functions after inflation is challenging to determine (the mode functions during a de Sitter phase are given by Hankel functions). Moreover, dimensional regularization requires that renormalization conditions be imposed to fix counter-terms, which typically give rise to logarithmic dependence on the renormalization scale (see, e.g., the appendix of dr-kaya for a discussion of how dimensional regularization works for the two-point function in de Sitter space). Yet, the Bunch-Davies mode functions exhibit the same UV behavior as their flat-space counterparts, and on dimensional grounds, one therefore expects $\phi^{*}\phi\right>\sim m^{2}$.
An alternative approach is the adiabatic regularization ad1 as applied to the two point function ad2 , where one subtracts the adiabatic mode contributions from the Bunch-Davies modes (note that this method is more suited to cosmological settings and does not require any renormalization conditions to be imposed). The evolution of Bunch-Davies mode function after inflation depends on its comoving wave-number $k$. For instance, modes that satisfy $k/a>\textrm{max}(m,H)$ at all times evolve adiabatically, and their contribution to the two-point function is canceled by adiabatic subtractions. One can then see that, since $m\gg H$ during the period of interest, adiabatic regularization also yields $\phi^{*}\phi\right>\sim m^{2}$.
As a result for $m\sim$ TeV, which is a reasonably small mass scale, one finds $\mu\gg 1$ unless $e$ is extremely small (the current bound on $\mu$ is $|\mu|<9\times 10^{-5}$ from COBE/FIRAS and proposed experiments like PIXIE aim to to push this bound to $|\mu|\sim 10^{-8}$).
A large $\mu$ appears from the same root cause as the cosmological constant problem, i.e. there is a fundamental lack of our understanding about the cosmological impact of coincident vacuum expectation values. One can try to estimate $\phi^{*}\phi\right>$ by applying the same amount of mismatch encountered in the cosmological constant problem. The naive dimensional or adiabatic regularization gives $\rho\right>\sim m^{4}$, where $\rho$ is the energy density, and assuming this must have the same order of magnitude as the dark energy today gives the rough mismatch factor $H_{0}^{2}M_{p}^{2}/m^{4}$, where $H_{0}$ is the current Hubble value. Because $\rho$ has mass dimension 4 and $\phi^{*}\phi$ has mass dimension 2, the mismatch for $\phi^{*}\phi\right>_{0}$ can be $H_{0}M_{p}/m^{2}$, which implies $\phi^{*}\phi\right>_{0}\sim H_{0}M_{p}$. Taking $\omega_{0}\sim T_{0}\sim 10^{-4}eV$, the second term in (26) gives $\mu\sim 10^{-5}e^{2}$, which can be within the reach of the upcoming measurements if $e\sim 1/10$.